### v1

# Advancing MAPF towards the Real World: A Scalable Multi-Agent Realistic Testbed (SMART)

Abstract
We present Scalable Multi-Agent Realistic Testbed (SMART), a realistic and efficient software tool for evaluating Multi-Agent Path Finding (MAPF) algorithms. MAPF focuses on planning collision-free paths for a group of agents. While state-of-the-art MAPF algorithms can plan paths for hundreds of robots in seconds, they often rely on simplified robot models, making their real-world performance unclear. Researchers typically lack access to hundreds of physical robots in laboratory settings to evaluate the algorithms. Meanwhile, industrial professionals who lack expertise in MAPF require an easy-to-use simulator to efficiently test and understand the performance of MAPF algorithms in their specific settings. SMART fills this gap with several advantages: (1) SMART uses a physics-engine-based simulator to create realistic simulation environments, accounting for complex real-world factors such as robot kinodynamics and execution uncertainties, (2) SMART uses an execution monitor framework based on the Action Dependency Graph, facilitating seamless integration with various MAPF algorithms and robot models, and (3) SMART scales to thousands of robots. In addition, we use SMART to explore and demonstrate research questions about the execution of MAPF algorithms in real-world scenarios. The code is publicly available at https://jingtianyan.github.io/publication/2025-smart.
1 Introduction
Figure 1: (a) A simulation environment in SMART. (b) The original 2-D grid map from the MovingAI benchmark. (c) Robot model used in the environment. (d) A 3D view of the environment.
The Multi-Agent Path Finding (MAPF) problem Stern et al. (2019) focuses on planning collision-free paths for large groups of robots within a known environment. State-of-the-art MAPF algorithms can plan paths for hundreds of robots within seconds Ma et al. (2019); Li et al. (2021c); Zhang et al. (2022); Li et al. (2022). However, these algorithms make several simplifying assumptions. First, they often rely on simplified robot models that ignore kinodynamic constraints while planning the robots’ paths. Second, they assume that robots can execute these paths perfectly, without accounting for uncertainties introduced by real-world factors. Third, they typically plan and evaluate paths on unrealistic discretized 2-D grid maps Stern et al. (2019). Therefore, while researchers have shown that MAPF algorithms have a wide range of applications Ma et al. (2017); Hönig et al. (2019); Ho et al. (2022); Li et al. (2023); Zhang et al. (2023b, c, 2024); Zang et al. (2025), the algorithms are mostly evaluated under simplified settings, leaving their performance in more realistic settings unclear.
MAPF researchers in both academia and industry require an efficient, realistic, and scalable software tool to evaluate MAPF algorithms in realistic settings. In academia, deploying hundreds of physical robots in a large environment to evaluate MAPF algorithms is impractical. Meanwhile, developing software that can evaluate large groups of robots is a non-trivial task for MAPF researchers. Researchers must create high-fidelity simulation environments and robot models to capture realistic factors, as well as develop essential software components such as controllers and execution frameworks to execute their plans. Moreover, since many researchers come from the AI community, learning such simulation tools can be time-consuming and require significant effort. While some software tools Schaefer et al. (2023); Heuer et al. (2024) have been developed to evaluate MAPF algorithms in realistic settings, they do not scale to hundreds of robots and require significant engineering efforts to integrate new MAPF algorithms and maps. In addition, industrial professionals possess realistic simulators for their applications of interest. However, they may not have the resources and expertise to implement state-of-the-art MAPF algorithms in their simulators. By having a realistic simulator that can easily interface with existing implementations of MAPF algorithms, industrial professionals can more easily understand and select a MAPF algorithm that fits their demands.
In this paper, we present SMART, a Scalable Multi-Agent Realistic Testbed, a software tool used to evaluate MAPF algorithms in realistic scenarios. SMART comprises three main components: (1) a physics-engine-based simulator based on Argos 3 Pinciroli et al. (2012), (2) an execution monitoring (EM) server, and (3) robot-specific executors. The simulator generates realistic simulation environments based on user-defined configurations. Fig. 1 shows an example environment in SMART created from a 2-D grid map from the MovingAI benchmark Stern et al. (2019). The EM server parses the user-provided MAPF paths, tracks the execution progress, and communicates with the executors. It leverages the Action Dependency Graph (ADG) Hönig et al. (2019), a MAPF execution framework, to ensure robust execution of the paths generated using various simplified robot models. Each robot is assigned an individual executor that receives and executes actions from the server. SMART incorporates key real-world factors, including robot kinodynamics, collision dynamics, communication delays, execution imperfections, and spatial inaccuracies caused by controller limitations.
Our contributions in this paper are: (1) We present SMART, an open-source software tool to evaluate MAPF algorithms in realistic scenarios, which can scale to thousands of robots while ensuring replicability, (2) we evaluate SMART with a variety of open-source implementations of MAPF algorithms and robot models, showing its flexibility, and (3) we show some important findings regarding the application of MAPF algorithms in realistic scenarios. We also use them as examples of how SMART can be used to address such problems.
Figure 2: (a) MAPF problem. (b) Associated MAPF plan. (c) ADG from the MAPF plan. Green ovals are actions, with the first element as the reference time from the plan, the second as the action type, and the last two as start and goal locations. Horizontal edges are Type-1, and vertical edges are Type-2. All robots initially face north.
Figure 3: System overview of SMART.
2 Background
2.1 MAPF Problem and MAPF Plan
A MAPF instance Stern et al. (2019) consists of an undirected graph $\mathcal{G}_{M}=(\mathcal{V}_{M},\mathcal{E}_{M})$ and a set of $I$ robots =$\mathcal{R}=\{r_{1},\dots,r_{I}\}$ with their corresponding start and goal locations. Given a MAPF instance, a MAPF algorithm aims at finding a MAPF plan, consisting of collision-free paths for all robots from their start to goal locations. Standard MAPF uses a simplified robot model by discretizing the time into timesteps and assuming that, at each timestep, every robot either moves to an adjacent vertex or stays at its current vertex. Two robots collide if they are at the same vertex or swap vertices at the same timestep. An example MAPF instance and plan on a 2-D grid graph is shown in Fig. 2 (a) and (b), respectively. The objective of MAPF is to minimize the sum-of-cost, defined as the sum of travel time.
Prior research has proposed many MAPF algorithms. Some methods focus on finding solutions with optimal or bounded-suboptimal guarantees Sharon et al. (2015); Barer et al. (2014); Li et al. (2021c), while others focus on scalability, solving MAPF problems involving thousands of robots Okumura et al. (2022); Okumura (2023). Anytime algorithms quickly find an initial solution and then iteratively refine it as time progresses, balancing solution quality with computational efficiency Li et al. (2021a); Huang et al. (2022). Additionally, some methods incorporate complex real-world factors, such as kinodynamics and unexpected delays, into the planning process Cohen et al. (2019); Atzmon et al. (2020); Yan and Li (2024a, b). They usually employ a two-level planner, where the high level employs MAPF-based methods to resolve collisions among agents, while the low level plans for each agent, consider real-world factors.
2.2 Executing MAPF Plans
Given a MAPF plan, robots can remain collision-free if they strictly follow it spatially and temporally. While existing path-tracking techniques allow for reasonable spatial accuracy, precise temporal adherence is more challenging Čáp et al. (2016), especially since MAPF plans are typically generated using imperfect robot models. Thus, we need techniques to execute MAPF plans safely given the potential delays. One naive way is synchronized execution, where agents need to wait for all robots to finish their current action before starting the next. This ensures strict compliance with the MAPF plan but significantly reduces efficiency. Alternatively, one can replan whenever a robot is delayed, but this is computationally heavy and thus does not scale well. Therefore, executing MAPF plans using an execution framework is preferred Čáp et al. (2016). The state-of-the-art execution framework is Action Dependency Graph (ADG) Hönig et al. (2019), which enforces passing orders defined by MAPF plans while accounting for delays. ADG can execute plans generated by MAPF planners of various simplified robot models Varambally et al. (2022).
An ADG, as shown in Fig. 2 (c), is a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E}_{1},\mathcal{E}_{2})$ that encodes dependencies between actions from a MAPF plan. The vertex set $\mathcal{V}=\{v_{i}^{k}:k\in[1,n^{i}],r_{i}\in\mathcal{R}\}$ represents all actions performed by each robot $r_{i}$, where each vertex $v_{i}^{k}$ corresponds to the $k$-th action of $r_{i}$, and $n^{i}$ is the number of actions of $r_{i}$. The edge set $\mathcal{E}_{1}$ includes Type-1 edges, which enforces the sequential execution of actions for each robot. For a robot $r_{i}$, a Type-1 edge $(v_{i}^{k},v_{i}^{k+1})$ ensures that action $v_{i}^{k}$ is completed before action $v_{i}^{k+1}$ starts. The edge set $\mathcal{E}_{2}$ includes Type-2 edges, which enforces the inter-robot passing order at shared vertices. For any two vertices $v_{i}^{k}$ and $v_{j}^{s}$ from $r_{i}$ and $r_{j}$, a Type-2 edge $(v_{i}^{k},v_{j}^{s})$ ensures that $v_{i}^{k}$ must be completed before $v_{j}^{s}$ can start. During execution, each action is performed according to these dependencies. The detailed execution procedure is described in Section 3.2.
2.3 Evaluating MAPF Plans
To evaluate MAPF algorithms, researchers typically generate a plan for a given MAPF instance and then execute it. While many MAPF problem benchmark instances are proposed for discretized 2-D-grid maps Stern et al. (2019); Gebser et al. (2018); Sturtevant (2012); Mohanty et al. (2020); Qian et al. (2024), few software tools are provided to execute the MAPF plans in more realistic environments. Therefore, most prior works are evaluated using simplified robot models that assume no delays and discretized time and space.
Meanwhile, some software tools are presented to evaluate the MAPF plans in more realistic settings. MRP-Bench Schaefer et al. (2023) leverages a high-fidelity, physics-based simulator Gazebo Koenig and Howard (2004) to simulate multi-robot systems. It integrates low-level controllers to account for robots’ kinodynamic constraints and potential delays. REMROC Heuer et al. (2024) also uses Gazebo to realistically evaluate MAPF plans, with a particular focus on human-shared environments. However, they have two main limitations. First, they often require a lot of engineering effort to use user-defined maps and MAPF planners, limiting their flexibility for different scenarios. Second, their scalability is constrained by the high computational overhead of Gazebo, which models many factors less relevant to MAPF research. MRP-Bench, for example, does not scale to more than 10 robots. Additionally, many MAPF researchers from the AI community may not be familiar with tools like Gazebo, making installation and usage a challenging process.
3 SMART Architecture
Fig. 3 shows an overview of SMART. It comprises three main modules: a simulator, an execution monitoring (EM) server, and executors. All modules are implemented in C++. The simulator generates simulation environments based on the user-provided configurations. The EM server parses the MAPF plan and monitors the execution status of each action throughout the simulation. When an action is ready for execution, the server sends the action to the corresponding robot via the network. The executor of each robot receives actions from the server, executes them using a controller, and synchronizes the execution status with the server accordingly.
3.1 Simulator
[…]
Figure 4: (a) Cycle in the ADG causing all robots to be stalled. (b) Adding safe stops between vertices to make the ADG acyclic.
3.2 Execution Monitoring (EM) Server
The EM server parses the MAPF plan and encodes its passing order into an ADG. It then manages the execution of the MAPF plan and communicates with the executors.
Server Initialization
During the initialization, we use the MAPF plan parser to construct an ADG. We pre-define the format to store the MAPF plan as a sequence of timed locations for each robot. The parser can take as input any valid MAPF plans of this format, regardless of the MAPF planners that generate the plans. This allows SMART to work with plan using varying simplified robot models. The parser first converts the path of each robot into a sequence of actions. Specifically, the parser infers the orientation of a robot by comparing consecutive vertices along its path. If the required orientation differs from the current orientation, a rotate action is added. The ADG is constructed based on the sequence of actions. ADG requires the MAPF plan to be 1-robust, ensuring that robots do not collide even if delayed by one timestep. However, ensuring a 1-robust plan is not standard in most MAPF models. As a result, if we directly build an ADG based on a non-1-robust plan, we may create deadlocks, such as those illustrated in Fig. 4 (a). While these robots are planned to rotate simultaneously in a cycle in the non-1-robust plan, with ADG, they will end up waiting indefinitely for others to complete their next actions. To accommodate a broader range of planners, motivated by Hönig et al. (2016), we add an extra vertex between consecutive vertices in SMART as shown in Fig. 4 (b), making the ADG compatible with these MAPF models.
Execution Managing
During execution, an execution monitor framework manages the progress of execution and communicates with the executors. This framework follows the method described in ADG Hönig et al. (2019) to track and manage execution. Each vertex in the ADG represents an action and can have one of three statuses: staged, enqueued, or finished. At first, all vertices are staged. A vertex transitions to the enqueued if it has no preceding vertices or (1) its preceding vertex connected by a Type 1 edge is either enqueued or finished, and (2) its preceding vertices connected by Type 2 edges are finished. Actions associated with enqueued vertices are then sent to the executors for execution, while other vertices remain staged. A vertex is marked as finished when the executor confirms the completion of its associated action with the server. Communication with the executors is managed by an RPC server built with rpclib [^1] Available at http://rpclib.net..
3.3 Executor
Each robot is launched with its own individual executor. As shown in Fig. 3, each executor consists of an RPC client, an action queue, and a controller. The RPC client establishes communication with the server, receiving enqueued actions and sending confirmation for finished actions. The action queue keeps track of the enqueued actions, which are executed in a first-in-first-out manner. A Proportional Integral Derivative (PID) controller is used to execute these actions by sending control commands to the simulator. Once an action is finished, a confirmation message is sent to the server. To allow robots to operate at higher speeds, consecutive actions of the same type are merged into one action in the queue.
3.4 Performance Metrics
SMART provides several built-in performance metrics, including (1) Average execution time (AET): The sum of the simulation times required by all robots to complete their assigned paths, divided by the number of robots, (2) Maximum execution time: The longest simulation time taken by any single robot to complete its path, and (3) Robot state over time: The state of an individual robot at different times, where the state of a robot includes the size of its action queue, its current position, and control output. Additionally, the open-source nature of our code allows researchers to easily derive custom metrics from SMART as needed. SMART also provides visualization for the MAPF plan executions, allowing researchers to better understand the behaviors of the robots.
4 System Evaluation
We test SMART on two machines: (1) a server with a 64-core AMD 7980X processor and 256 GB of memory, and (2) a laptop with a 4-core Intel i7-6700HQ processor and 16 GB of memory. Both machines are equipped with Ubuntu 20.04. We use machine (1) to evaluate the scalability and applicability of SMART in Sections 4.1 and 4.2. We use both machines to evaluate CPU and memory usage in Section 4.3.
4.1 Scalability and Runtime Performance
This section evaluates the scalability and runtime performance of SMART by analyzing its simulation speed, which is defined as the ratio of the total number of simulation updates to the real-world time (in seconds). A simulation update refers to the periodic recalculation of the status of robots and the environment based on the input commands. In all experiments, the simulation update period is set to 0.1 simulation seconds. A higher simulation speed shows that more simulation updates can be finished within the same unit of time, indicating better scalability. For example, if the simulation speed is 100 and the update period is 0.1 simulation seconds, the simulation world runs 10 times faster than real time. We assess scalability on two environments converted from: warehouse (warehouse-10-20-10-2-1, size: 161$\times$63) from the MovingAI benchmark Stern et al. (2019) and sortation-center (size: $500\times 140$) from the League of Robot Runner MAPF competition Chan et al. (2024). MAPF plans are generated using MAPF-LNS2222Code at https://github.com/Jiaoyang-Li/MAPF-LNS2 Li et al. (2022), a state-of-the-art MAPF planner, with a progressively increasing number of robots and a time limit of 60 seconds. These plans are then executed in SMART without visualization. Notably, the MAPF instances from warehouse support a maximum of 1,000 robots, whereas MAPF-LNS2 can handle up to 2,000 robots within the time limit on sortation-center.
Figure 5: Simulation speed on warehouse and sortation-center maps.
As shown in Fig. 5, SMART supports simulations with up to 2,000 robots. The more robots and the larger the map, the smaller the simulation speed. Nevertheless, SMART maintains a simulation speed of around 100 with 1,000 robots in warehouse. Even in the larger sortation-center, the simulation speed is consistently greater than 10.
Robots (#) average execution time maximum execution time
100 263.36 $\pm$ 0.01 469.80 $\pm$ 0.00
400 385.76 $\pm$ 0.02 737.60 $\pm$ 0.42
1000 612.04 $\pm$ 0.07 1389.43 $\pm$ 0.83
Table 1: Replicability using different numbers of robots.
4.2 Replicability
We evaluate the consistency of SMART by examining whether the same MAPF plans yield the same outcomes across multiple runs. Due to the realistic setting we used, uncertainty may arise from factors such as network communication delays or controller variability. Using the same warehouse map and MAPF plans as in the previous section, we execute SMART 10 times for each MAPF plan. For each run, we record both the average execution time and the maximum execution time. As shown in Table 1, SMART demonstrates strong replicability across runs, with minor execution time variations. While these variations increase slightly with more robots, they remain minimal even for up to a thousand robots.
4.3 CPU and Memory Usage
We evaluate the CPU and memory usage of SMART on both machines (1) and (2). We use machine (2) to demonstrate that SMART can run on resource-constrained machines. To evaluate usage, we generate MAPF plans with MAPF-LNS2 and run SMART on the warehouse map with different numbers of robots, each with 5 random MAPF instances selected from the MAPF benchmark Stern et al. (2019). All simulations are parallelized on all CPU cores on both machines. We record the peak CPU and memory usage while running the simulation and compute the average of all instances for each number of robots. Table 2 shows the results. We observe that SMART has reasonable CPU usage and low memory usage on both machines, making it possible to run on low-spec personal laptops. In machine (2), the simulation speed is slower than in machine (1) due to an inferior processor. For reference, we run the Gazebo-based simulator Varambally et al. (2022) on machine (2) with 50 robots in an environment of size 31$\times$12. Since Gazebo operates on a single thread, it utilizes 100% of a single CPU core, consumes 2.09 GB of memory, and achieves a simulation speed of 5.40.
Machine Robots (#) CPU Usage (%) Memory Usage (GB) Simulation Speed
(1) 100 16.48 $\pm$ 0.58 0.19 $\pm$ 0.26 334.90 $\pm$ 1.54
1000 46.02 $\pm$ 0.98 0.65 $\pm$ 0.03 73.71 $\pm$ 0.40
(2) 100 27.96 $\pm$ 0.39 0.04 $\pm$ 0.00 143.74 $\pm$ 1.03
1000 59.64 $\pm$ 0.32 0.63 $\pm$ 0.02 19.45 $\pm$ 0.09
Table 2: Comparison of CPU and Memory Usage
5 Problems from MAPF to Realistic Scenarios
We argue that SMART has several broader impacts on the MAPF research community. First, it allows researchers to evaluate MAPF plans from arbitrary MAPF planners in a simulator with realistic robot kinodynamics models and execution delays caused by real-world factors. Second, it bridges the gap between MAPF planning algorithms and executing the planned paths in the real world, bringing researchers more guidance on designing high-quality MAPF algorithms for the real world. Third, SMART can be used to study research problems regarding executing MAPF plans in the real world.
In this section, we study research problems focusing on the gap between MAPF planning and real-world execution of the planned paths. We study these problems using six simulation environments converted from six maps from the MovingAI benchmark Stern et al. (2019): empty (empty-32-32, size: 32$\times$32), random (random-64-64-10, size: 64$\times$64), room (room-64-64-16, size: 64$\times$64), den312d (den312d, size: 65$\times$81), maze (maze-32-32-4, size: 32$\times$32), and warehouse (warehouse-10-20-10-2-1, size: 161$\times$63).
Feature All Features SOC Rotations (#) Type-1 edges (#) Type-2 edges (#) Conflict robot pairs (#) Robots (#)
MAPE 0.0342 0.1778 0.1793 0.2414 0.2826 0.3175 0.4317
Table 3: Quadratic Regression Results. Except for the first column, each column means that the model is trained with only a single feature
5.1 Impact of Optimality
We first use SMART to study the practical effectiveness of the objective used by MAPF algorithms. One of the most widely used objectives in MAPF research is the Sum of Costs (SOC), which measures the total number of timesteps for all robots to reach their goals. While extensive efforts have been made to develop algorithms that optimize SOC optimally or sub-optimally, it is unclear, for example, whether reducing SOC by 1% translates to a similar degree of improvement in AET in a more realistic simulation. In this section, we use SMART to assess the correlation between improvements in SOC and AET. Moreover, we want to explore better objectives than SOC in terms of evaluating the quality of the MAPF plan in more realistic settings.
5.1.1 Experiment Setup
We generate MAPF plans with varying SOC for a specific MAPF instance by using MAPF-LNS [^3] Code at https://github.com/Jiaoyang-Li/MAPF-LNS Li et al. (2021a), an anytime method that iteratively refines MAPF plans to improve SOC within a predefined runtime limit. For each map, we select three different numbers of robots and five sets of different start and goal locations for robots. For each MAPF instance, we record all plans generated in each iteration of MAPF-LNS within the runtime limit of 60 seconds. We then execute these plans in SMART.
Figure 6: Relationship between SOC and AET. Each dot represents a MAPF plan and its execution result.
5.1.2 Results and Analysis
We first study whether SOC can measure the real-world execution time of a MAPF plan. As shown in Fig. 6, SOC and AET show a strong positive correlation across all environments and all numbers of robots, indicating that SOC captures the trend of AET. However, the correlation is not perfectly monotonic. For the same SOC value, there are noticeable variations in AET, suggesting that SOC alone cannot account for all factors influencing execution time. This highlights the limitation of SOC as the sole objective of the MAPF problem.
To explore other factors that influence execution time, we extract additional features from the ADG, including (1) Type-2 edges (#), the number of type-2 edges, (2) Type-1 edges (#), the number of type-1 edges, (3) Rotations (#), the number of rotation actions, (4) Conflict robot pairs (#), the number of robot pairs connected by type-2 edges, reflecting dependencies in their passing orders, and (5) Robots (#), the number of robots. We analyze the correlation between these features and AET by using a quadratic regression model to predict AET based on these features. We measure prediction accuracy by using the mean absolute percentage error (MAPE). As shown in Table 3, SOC exhibits the highest predictive power among individual features, while other features, such as the number of Type-2 edges, Type-1 edges, and Rotations, also show correlation with AET. Meanwhile, the model incorporating all features achieves a much lower MAPE than any single feature, indicating that multiple features influence the execution time. In conclusion, incorporating features in addition to SOC is crucial to developing a more accurate objective function for execution time estimation during MAPF planning.
Figure 7: AET comparison for MAPF with different robot models.
5.2 Impact of Different Robot Models
The standard MAPF problem relies on a simplified model that ignores critical execution factors such as kinodynamics and potential delays. This simplification creates a gap between planning and real-world execution. Thus, recent research has introduced more realistic MAPF models that consider real-world factors. An ideal MAPF model should capture the essential aspects of robot kinodynamics while omitting unnecessary details. As SMART supports various robot models, we conduct an empirical comparison of different MAPF models to identify the most effective one in real-world scenarios. In Varambally et al. (2022), a similar problem is studied, but the experiments are limited to a warehouse map and consider no more than 50 robots. Additionally, we compare a wider range of robot models.
5.2.1 Experiment Setup
We evaluate the impact of incorporating robot kinodynamics and potential delays. For the kinodynamics models, we consider: (1) the standard MAPF model, (2) the MAPF model with rotation, and (3) the MAPF model that accounts for the full kinodynamics of robots, including rotation, speed, and acceleration. To model delays, we use the $k$-robust delay model Atzmon et al. (2020), which ensures that robots can avoid collisions even if delayed by up to $k$ timesteps. By combining kinodynamics and delay factors, we evaluate the following MAPF algorithms: (1) MAPF, $k=i$: This method uses the standard MAPF model and a $k$-robust delay model of $k=i$. (2) MAPF+rotation, $k=i$: This extends the standard MAPF model by incorporating in-place rotation but does not account for other kinematics. It also includes a $k$-robust delay model. (3) MAPF+kinodynamics: Derived from Yan and Li (2024a), this method incorporates accurate kinodynamic constraints of robots. We use PBS Ma et al. (2019) to handle conflicts between robots under these models with a runtime limit of 60 seconds. We run these methods to generate MAPF plans for all maps using 25 “random scenarios” with a progressive number of robots until the success rate is smaller than 20%. Finally, these plans are executed using SMART.
Figure 8: Paths generated on different maps.
5.2.2 Results and Analysis
As shown in Fig. 7, the method incorporating the full kinodynamics of robots achieves the shortest AET among all the models. Additionally, the planner considering robot rotations demonstrates much better AET than the standard MAPF model. These results indicate that incorporating more accurate kinodynamic models into planning can improve execution time. However, we also observe trade-offs associated with these models. More complex kinodynamic models lead to longer runtimes during planning. Meanwhile, the $k$-robust model achieves better or comparable performance with fewer robots with larger values of $k$; however, its performance declines significantly as the number of robots increases, resulting in lower solution quality and reduced scalability. Interestingly, in warehouse environments, MAPF+kinodynamics does not provide significant improvements over the MAPF+rotation. To better understand this, we visualize the MAPF plans in Fig. 8, comparing the warehouse map with random and room maps. The planned paths show that MAPF+rotation and MAPF+kinodynamics generate paths with more consecutive movements and fewer conflict vertices in warehouse maps than in the other two maps. We hypothesize that this minimizes the potential benefits of accounting for kinodynamics in such environments. In summary, while more accurate kinodynamic models generally improve execution time in realistic settings, their increased computational complexity can limit scalability. Additionally, the $k$-robust model is less efficient and may even harm performance, particularly in more congested scenarios.
Figure 9: AET using different MAPF models and planners.
5.3 Optimality and Robot Model
From Sections 5.1 and 5.2, we show that both improving SOC and incorporating a more accurate kinodynamic model during planning can lead to better AET during execution. However, both of them require additional computation time. Considering the limited planning time available in real-world applications, this section explores the balance between these factors to provide practical insights.
5.3.1 Experiment Setup
We generate MAPF plans with varying SOC for three robot models: (1) standard MAPF, (2) MAPF with rotation, and (3) MAPF with kinodynamics. To obtain optimal solutions, we use CBS [^4] Code at https://github.com/Jiaoyang-Li/CBSH2-RTC Li et al. (2021b) for the standard MAPF model and CBS with rotation [^5] Code at https://github.com/YueZhang-studyuse/MAPF_T Zhang et al. (2023a) for the rotation model. For suboptimal plans, we combine PBS with all three models, as described in Section 5.2. To analyze the relationship between AET and SOC, we create a naive anytime planner by running PP with random restarts and recording the MAPF plan with the lowest SOC until reaching a runtime limit of 60 seconds. We do not use existing anytime solvers, such as MAPF-LNS, because, to our best knowledge, no available implementation considers robot models with kinodynamic constraints.
5.3.2 Results and Analysis
As shown in Fig. 9, planners using more accurate robot models consistently achieve better AET across all maps, even when the methods using less accurate models are optimal. This indicates that incorporating more accurate kinodynamic robot models into planning may be more important than obtaining optimal solutions with less accurate models. Notably, as discussed in Section 5.1, since SOC is not a perfect metric, the optimal planner does not always achieve the best AET. Meanwhile, execution frameworks may also become a bottleneck. Since ADG adopts an overly conservative strategy, assuming robots can be delayed indefinitely, it compromises the solution quality of the MAPF plan in certain cases. In conclusion, the limiting factors for achieving better AET lie in the more accurate robot models or the execution frameworks with better performance.
6 Conclusion
We introduced SMART, a scalable testbed for evaluating MAPF algorithms under realistic settings. SMART bridges the gap between methods that use simplified models and their real-world deployments. SMART is compatible with existing 2-D grid-based benchmarks, making it straightforward for researchers to evaluate their methods in more realistic environments. Our experiments demonstrated that SMART can handle thousands of robots while maintaining high simulation speeds. We highlight several key challenges arising from the application of MAPF in realistic scenarios. We further show how SMART can be used to address these challenges. Future work includes extending the framework to address lifelong MAPF scenarios, integrating it with other execution frameworks Coskun and O’Kane (2019); Berndt et al. (2023); Feng et al. (2024), and interfacing it with real robots.
Acknowledgments
The research was supported by the National Science Foundation (NSF) under grant number #2328671 and a gift from Amazon. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the sponsoring organizations, agencies, or the U.S. government.